OpenStudy (anonymous):

I can't figure out the equivalent equation of 2*-4=18

OpenStudy (zzr0ck3r):

there is nothing to solve you wrote -8 = 18 which is never true

OpenStudy (anonymous):

im not catching what u mean?

OpenStudy (zzr0ck3r):

uri?

OpenStudy (uri):

your question doesn't make sense @codilee98

OpenStudy (anonymous):

Ill write the numbers. 2*-4=18

OpenStudy (zzr0ck3r):

are you sure there is not a x someplace?

OpenStudy (uri):

lool @zzr0ck3r funny cx.

OpenStudy (anonymous):

2x-4=18

OpenStudy (zzr0ck3r):

2 times -4 is equal to -8 this is never equal to 18

OpenStudy (uri):

2x-4=18 2x and 2* are different here. lmao @codilee98

OpenStudy (uri):

add 4 both sides @codilee98

OpenStudy (zzr0ck3r):

This has by far been the best 4 minutes I have ever spent on this site

OpenStudy (anonymous):

i did add 4 to both sides then i got stuck on how to finish

OpenStudy (fibonaccichick666):

what do you get when you add 4 to both sides?

OpenStudy (anonymous):

i get 8

OpenStudy (fibonaccichick666):

I have been following this amazing 4 minutes zz lol

OpenStudy (uri):

adding 4 both sides gives you 2x=22

OpenStudy (fibonaccichick666):

ok, so you can't just get 8 when you add four to both sides of an equation, can you show me how you got 8?

OpenStudy (uri):

2x-4=18 +4 +4 2x=22 @codilee98 get it?

OpenStudy (anonymous):

+4 on both sides

OpenStudy (uri):

-4 +4 = cancelled.

OpenStudy (fibonaccichick666):

how like show that, because no matter what, you should still have an = sign and an x somewhere

OpenStudy (anonymous):

then beside that it shows 3+x/5=5

OpenStudy (uri):

2x=22 divide 22 by 2 22/2 =? @codilee98

OpenStudy (uri):

whatever you did was wrong @codilee98 forget it,hon.

OpenStudy (anonymous):

3+x/5=5

OpenStudy (uri):

whats this now?

OpenStudy (anonymous):

this was the other numbers beside this problem

OpenStudy (fibonaccichick666):

can you type the whole question for us?

OpenStudy (uri):

this is the other ques?

OpenStudy (anonymous):

the numbers i just typed out i need to find the equivalent equation of 3+x/5=5

OpenStudy (uri):

this is the other question right?